# Stability Properties of Multi-Order Fractional Differential Systems in 3D

### v1

@@ 3.1 Homogeneous linear systems
From eq. (3), we can see that $Q$ is continuous on $\mathbb{R}$, that $\lim_{s\to\infty}Q(s)=\infty$, and that $Q(0)=-\det A$. Hence, if $\det A>0$ then $Q(0)<0$ and thus $Q$ must have a real zero in the interval $(0,\infty)$. Therefore, the statement of Theorem 1 asserts the instability in case (a). We will give the proof of (b) at the end of Subsection 3.1. ∎
@@
From $d<0$ we see that $Q(0)\neq 0$. Moreover, since $Q(\bar{s})=\overline{Q(s)}$, it follows that $Q(\omega\mathrm{i})=0$ with some $\omega>0$ if and only if $Q(-\omega\mathrm{i})=0$. Thus, the statement follows directly from the definitions (6) and (7) of $h_{1}$ and $h_{2}$. ∎
@@
where $0<\varepsilon<R$ are chosen such that all zeros of $Q$ in the open right half of $\mathbb{C}$ are located inside of the open region $C_{\varepsilon,R}\subset\mathbb{C}$ whose boundary is formed by these four curves. This is possible if we choose $\epsilon$ sufficiently small (since $Q(0)>0$, the continuity of $Q$ implies that $Q$ has no zeros $s$ with $|s|\leq\varepsilon$) and $R$ sufficiently large (because $\lim_{|s|\to\infty}|Q(s)|=\infty$ uniformly in $\arg s$). As in Diethelm et al. (2022, pp. 1332–1333), we see that the number of zeros of $Q$ in the open right half of $\mathbb{C}$ is equal to the number of zeros in $C_{\varepsilon,R}$ which, since $Q$ is analytic in $C_{\varepsilon,R}$, is equal (by the argument principle) to the number $Z$ of times that the curve $Q(\cup_{k=1}^{4}\gamma_{k})$ encircles the origin in the positive direction. If $Z>0$ then this curve must encircle the origin at least once, and so the curve must cross the positive and negative real and imaginary axes. Thus, in particular there is some $s^{*}\in\cup_{k=1}^{4}\gamma_{k}$ such that $Q(s^{*})\in(-\infty,0)$.
@@
    $d\leq-b(b\tilde{\rho}_{2}+c\tilde{\rho_{1}})^{\frac{\beta_{2}}{\beta_{4}-\beta_{1}}}\rho_{2}-c(b\tilde{\rho}_{2}+c\tilde{\rho_{1}})^{\frac{\beta_{1}}{\beta_{4}-\beta_{1}}}\rho_{1};$
@@
    $d\leq-a(a\tilde{\rho}_{3}+c\tilde{\rho_{1}})^{\frac{\beta_{3}}{\beta_{4}-\beta_{1}}}\rho_{3}-c(a\tilde{\rho}_{3}+c\tilde{\rho_{1}})^{\frac{\beta_{1}}{\beta_{4}-\beta_{1}}}\rho_{1};$
@@
    $d\leq-a(a\tilde{\rho}_{3}+b\tilde{\rho_{2}})^{\frac{\beta_{3}}{\beta_{4}-\beta_{2}}}\rho_{3}-b(a\tilde{\rho}_{3}+b\tilde{\rho_{2}})^{\frac{\beta_{2}}{\beta_{4}-\beta_{2}}}\rho_{2}.$
@@
$h_{2}(\omega)$ $>$ $\omega^{\beta_{4}}\sin\frac{\beta_{4}\pi}{2}-\left(b\sin\frac{\beta_{2}\pi}{2}+c\sin\frac{\beta_{1}\pi}{2}\right)\omega^{\beta_{1}}$
$=$ $\omega^{\beta_{1}}\sin\frac{\beta_{4}\pi}{2}\left(\omega^{\beta_{4}-\beta_{1}}-\left(b\tilde{\rho}_{2}+c\tilde{\rho}_{1}\right)\right)$
for every $\omega>1$, and thus $h_{2}((b\tilde{\rho_{2}}+c\tilde{\rho_{1}})^{1/(\beta_{4}-\beta_{1})})>0$. This implies that $1<\omega_{0}<(b\tilde{\rho_{2}}+c\tilde{\rho_{1}})^{1/(\beta_{4}-\beta_{1})}$. Hence, in view of eq. (9) and our assumption on the value of $d$,
@@
$>$ $-b(b\tilde{\rho_{2}}+c\tilde{\rho_{1}})^{\beta_{2}/(\beta_{4}-\beta_{1})}\rho_{2}$
${}-c(b\tilde{\rho_{2}}+c\tilde{\rho_{1}})^{\beta_{1}/(\beta_{4}-\beta_{1})}\rho_{1}-d$
@@
which completes the proof in case (i). For cases (ii) and (iii), we proceed in the same way. ∎

### v2

@@ 3.1 Homogeneous linear systems
From eq. (3), we can see that $Q$ is continuous on $\mathbb{R}$, that $\lim_{s\to\infty}Q(s)=\infty$, and that $Q(0)=-\det A$.
Hence, if $\det A>0$ then $Q(0)<0$ and thus $Q$ must have a real zero in the interval $(0,\infty)$. Therefore, the statement of Theorem 1 asserts the instability in case (a).
We will give the proof of (b) at the end of Subsection 3.1. ∎
@@
Since $Q(\bar{s})=\overline{Q(s)}$, it follows that $Q(\omega\mathrm{i})=0$ with some $\omega>0$ if and only if $Q(-\omega\mathrm{i})=0$. Moreover, from $d<0$ we see that $Q(0)\neq 0$. Thus, the statement follows directly from the definitions (6) and (7) of $h_{1}$ and $h_{2}$. ∎
@@
where $0<\varepsilon<R$ are chosen such that all zeros of $Q$ in the open right half of $\mathbb{C}$ are located inside of the open region $C_{\varepsilon,R}\subset\mathbb{C}$ whose boundary is formed by these four curves. This is possible if we choose $\varepsilon$ sufficiently small (since $Q(0)>0$, the continuity of $Q$ implies that $Q$ has no zeros $s$ with $|s|\leq\varepsilon$) and $R$ sufficiently large (because $\lim_{|s|\to\infty}|Q(s)|=\infty$ uniformly in $\arg s$). As in Diethelm et al. (2022, pp. 1332–1333), we see that the number of zeros of $Q$ in the open right half of $\mathbb{C}$ is equal to the number of zeros in $C_{\varepsilon,R}$ which, since $Q$ is analytic in $C_{\varepsilon,R}$, is equal (by the argument principle) to the number $Z$ of times that the curve $Q(\cup_{k=1}^{4}\gamma_{k})$ encircles the origin in the positive direction. If $Z>0$ then this curve must encircle the origin at least once, and so the curve must cross the positive and negative real and imaginary axes. Thus, in particular there is some $s^{*}\in\cup_{k=1}^{4}\gamma_{k}$ such that $Q(s^{*})\in(-\infty,0)$.
@@
    $d\leq-b(b\tilde{\rho}_{2}+c\tilde{\rho_{1}})^{\frac{\beta_{2}}{\beta_{4}-\beta_{2}}}\rho_{2}-c(b\tilde{\rho}_{2}+c\tilde{\rho_{1}})^{\frac{\beta_{1}}{\beta_{4}-\beta_{2}}}\rho_{1};$
@@
    $d\leq-a(a\tilde{\rho}_{3}+c\tilde{\rho_{1}})^{\frac{\beta_{3}}{\beta_{4}-\beta_{3}}}\rho_{3}-c(a\tilde{\rho}_{3}+c\tilde{\rho_{1}})^{\frac{\beta_{1}}{\beta_{4}-\beta_{3}}}\rho_{1};$
@@
    $d\leq-a(a\tilde{\rho}_{3}+b\tilde{\rho_{2}})^{\frac{\beta_{3}}{\beta_{4}-\beta_{3}}}\rho_{3}-b(a\tilde{\rho}_{3}+b\tilde{\rho_{2}})^{\frac{\beta_{2}}{\beta_{4}-\beta_{3}}}\rho_{2}.$
@@
$h_{2}(\omega)$ $>$ $\omega^{\beta_{4}}\sin\frac{\beta_{4}\pi}{2}-\left(b\sin\frac{\beta_{2}\pi}{2}+c\sin\frac{\beta_{1}\pi}{2}\right)\omega^{\beta_{2}}$
$=$ $\omega^{\beta_{2}}\sin\frac{\beta_{4}\pi}{2}\left(\omega^{\beta_{4}-\beta_{2}}-\left(b\tilde{\rho}_{2}+c\tilde{\rho}_{1}\right)\right)$
for every $\omega>1$, and thus $h_{2}((b\tilde{\rho_{2}}+c\tilde{\rho_{1}})^{1/(\beta_{4}-\beta_{2})})>0$. This implies that $1<\omega_{0}<(b\tilde{\rho_{2}}+c\tilde{\rho_{1}})^{1/(\beta_{4}-\beta_{2})}$. Hence, in view of eq. (9) and our assumption on the value of $d$,
@@
$>$ $-b(b\tilde{\rho_{2}}+c\tilde{\rho_{1}})^{\beta_{2}/(\beta_{4}-\beta_{2})}\rho_{2}$
${}-c(b\tilde{\rho_{2}}+c\tilde{\rho_{1}})^{\beta_{1}/(\beta_{4}-\beta_{2})}\rho_{1}-d$
@@
which completes the proof in case (i).
For cases (ii) and (iii), we proceed in the same way. ∎